OpenStudy (anonymous):

Check my work!.... Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. -x + 3y = 9 -x -x 3y = -x + 9 slope: -x y-intercept: 9 Is that right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

add +x on both sides

OpenStudy (anonymous):

Oh! YEAH!! I knew I forgot something... lol. Thanks for catching that... xD

OpenStudy (anonymous):

-x+x cancel eachother remainig 3y=x+9 now divede bothsides by 3 y=x/3+3 now compare with standard form

OpenStudy (anonymous):

slope=1/3 y intercept=3

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

i have a question for you?will you tell me whats the meaning of lol

OpenStudy (anonymous):

LOL: laugh out loud.

OpenStudy (anonymous):

thanks